Please help me solve this problem: 13-(-5m+3)=3+3m-17
Found 2 solutions by Theo, MathLover1:
Answer by Theo(13342)

the equation is:

13 - (-5m+3) = 3 + 3m-17

first remove parentheses.

- (-5m+3) is equivalent to -1 * (-5m+3) which is equal to (-1 * -5m) + (-1 * 3).

-1 * -5m is equal to 5m.

-1 * 3 is equal to -3.

- (-5m+3) is therefore equivalent to 5m - 3

your equation of 13 - (-5m+3) = 3 + 3m-17 becomes:

13 + 5m - 3 = 3 + 3m - 17

combine like terms to get:

10 + 5m = 3m - 14

subtract 3m from both sides of the equation and subtract 10 from both sides of the equation to get:

5m - 3m = -14 - 10

combine like terms to get:

2m = -24

divide both sides by 2 to get:

m = -12

that should be your solution if we did this correctly.

confirm by replacing m with -12 in the original equation of 13 - (-5m+3) = 3 + 3m-17 to get:

13 - (-5 * -12 + 3) = 3 + 3 * -12 - 17

simplify to get:

13 - (60 + 3) = 3 - 36 - 17

simplify further to get:

13 - 63 = 3 - 36 - 17

combine like terms to get:

-50 = -50

this confirms the solution is good.

the solution is m = -12.

Answer by MathLover1(20849)